OpenStudy (anonymous):

radical 3 plus radical one-third gives what? how to solve?

OpenStudy (saifoo.khan):

that cannot be solved further!

OpenStudy (anonymous):

well that is the problem that i am given to solve.. ??? confused!

OpenStudy (saifoo.khan):

\[\sqrt{3} + \sqrt{ \frac13}\] \[\sqrt{3} + \frac{\sqrt 1}{\sqrt 3}\] \[\sqrt{3} + \frac{ 1}{\sqrt 3}\]

OpenStudy (saifoo.khan):

Take the LCD = radcal 3 \[\frac{\sqrt3 ( \sqrt 3) + 1 }{\sqrt3}\]\[\frac{3 + 1}{\sqrt 3}\]Can you solve now?

OpenStudy (anonymous):

not rly i think lol

OpenStudy (saifoo.khan):

lol, add 3 + 1.. that's it

OpenStudy (anonymous):

wait so is the answer 4 over radical 3????

OpenStudy (saifoo.khan):

Yes ma'am.

OpenStudy (anonymous):

yea! my brain may have some smarticles once ppl explain stuff to me lol

OpenStudy (saifoo.khan):

Haha. please hit good answer. Tx.

OpenStudy (anonymous):

Mhmm! :D